Find the area bounded by y=cosx, the x-axis and between x=−π/2 to x=π/2. Draw a sketch for yourself to help you find the integral.
1
Expert's answer
2020-05-24T18:59:06-0400

The region bounded by "y=cosx" , "x"−axis and between "y = -\frac{\pi}{2}" and "y = \frac{\pi}{2}" is the region below the curve as shown in figure:



Hence Area of required region is

"A = \int_{-\pi/2}^{\pi/2} y dx = \int_{-\pi/2}^{\pi/2} cos(x)dx \\ =sin(x)|_{-\pi/2}^{\pi/2} = -sin(-pi/2) +sin(pi/2) = -(-1)+1 = 2"